OpenStudy (anonymous):

just making sure im doing this all right.. 7 over 13 - 5 over 13 equals 2 over 13 right? (fractions)

OpenStudy (anonymous):

your trying to say \[\frac{ 7 }{ 13 }-\frac{ 5 }{ 13 }=\frac{ 2 }{ 13 }\]

OpenStudy (anonymous):

you'er*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yea that right

OpenStudy (anonymous):

thanks. this is just starting to get so confusing. like the problem